OpenStudy (anonymous):

Write an equation in point slope form of the line that passes through (5, 0), and whose slope is -2. y = -2(x – 5) y – 5 = -2(x - 5) y = -2(x + 5) y = 2(x – 5)

OpenStudy (anonymous):

y=mx+c

OpenStudy (zepp):

You have the slope and a points, plug them in \(y=mx+c\)

OpenStudy (anonymous):

here m= -2, put (5,0) at the place of x and y and solve for c

OpenStudy (zepp):

Find c \(y=mx+c \\ 0=-2(5)+c\\ 0=-10+c\\ 0--10=c\\ c=10\)

OpenStudy (zepp):

\(y=-2x+10\) Factor -2 out. \(y=-2(x-5)\)